# Driving as a Diagnostic Tool: Scenario-based Cognitive Assessment in Older Drivers From Driving Video

### v1

@@ 3.2 Executing Video-based Model
To evaluate our video-based embedding framework, we conducted experiments across two distinct driving scenarios that capture both emergency maneuvers and straight-path driving behaviors. The aim was to analyze whether these movement types could be effectively distinguished through video-derived embeddings.
The experiments were conducted under the following evaluation protocols:
@@
    Random Sampling: Video segments are randomly split into training and testing sets without enforcing driver-level separation.
@@
    Driver-level Separation: A leave-$k$-drivers-out cross-validation strategy is employed, where $k=5$ drivers were excluded during training and used exclusively for testing. This protocol was repeated over $r=3$ independent runs with different left-out driver groups, and the final metrics were averaged to capture subject variability.
In our experiments, the parameters are set as: (1) Input image size (960x752x3), (2) Embedded dimension (6144x1), (3) Frame Rate—$\{1,10\}$, (3) Low-dimension size— $n\in\{50,100,200\}$, and (4) Distance Metrics—L2 and centroid-distance. These experiments provide a comprehensive understanding of how video-derived embeddings behave under varying sampling conditions and data splits and lay the foundation for downstream classification tasks.
@@ 4.1 Model Performance on Different Scenarios
Table 3 reveals the performance of the video-based framework trained to identify two cognitive aging-groups (Normal-aging vs. AD-aging) across two scenario categories—Freeway interchange and Interstate. The $\Delta$ row (without driver-level separation) indicates significant separation on all metrics $\mathbf{(a,P,R,F1)}$ and validates the selection of these two contrastive scenarios to classify aging groups. The performance difference (grey colored rows) between the two scenarios indicates that there are prominent vision features in these specific scenarios that can successfully distinguish the aging classes solely by looking at the driving video features.
@@
Furthermore, the driver-level separation yields slightly widened separation, which can be observed on all metrics $\mathbf{(a,P,R,F1)}$ in the $\Delta_{\text{ds}}$ row. It surpassed the performance difference achieved in the without Driver-level Separation case; for example, the accuracy difference went up from $\Delta$ = 15.91% to $\Delta_{\text{ds}}$ = 17.64%, indicating a 1.73% increment in performance difference. The $\Delta_{\text{ds}}$ aligns with achieving a significant performance difference on two contrastive scenarios to classify aging groups while there is no “Subject-level overlapping”. It also validates that the Fwy-int. scenario includes more prominent driving movement features compared to its counterpart. On the other hand, the Interstate scenario performs significantly worse, yielding lower accuracy and F1-score. Therefore, this observation emphasizes the Freeway interchange as a more behaviorally informative and cognitively sensitive driving scenario, where complex maneuvers, short distances and lower speed variability have more nuanced driver responses.
@@
We compared our results with a recent study [22] that used a multi-modal (demographics and sleep data) framework to detect cognitive decline in older drivers. Our method outperformed both without Driver-level Separation ([22] accuracy 68.64% vs. our Fwy-int. accuracy 71.03%) and with Driver-level Separation ([22] accuracy 70.48% vs. our Fwy-int. max accuracy 72.32%) cases on the RWRAD dataset. This highlights the effectiveness of scenario-specific vision features as non-invasive digital biomarkers for cognitive aging classification. Moreover, another recent study by Danda et al. [23] found that amyloid-positive older adults exhibited higher average speeds, intermittent speed variability and increased acceleration on freeway on-ramp driving (similar to Fwy-int. scenario) suggesting less adaptive driving. In clinical research, Amyloid is a protein that accumulates abnormally in the brain and forms plaques linked to the development of AD. Patients with cognitive impairment who test amyloid-positive are typically diagnosed as having AD, whereas amyloid-negative patients are generally excluded from an AD diagnosis [24, 25]. These findings confirm that Freeway ramp driving requires increased cognitive load involving behavioral reactions to rapidly appearing traffic and abrupt speed shifts, combining perception with motor control [23]. Additionally, they underscore our results of higher classification accuracy ($\mathbf{a}$ = 71.03%) and F1-score ($\mathbf{F1}$ = 0.6712) in Freeway interchange scenarios compared to the Interstate settings ($\mathbf{a}$ = 55.12%, $\mathbf{F1}$ = 0.5368) reported in Table 3. Although their approach combined driving patterns with physiological signals, it validates the relevance of using contrastive and challenging roadway scenarios in cognitive classification tasks. Therefore, it aligns with our findings that the Freeway interchange scenario provides more informative and strong behavioral cues that the vision model can leverage to better separate cognitive aging groups.
@@ 4.2 Percentage of Misclassification
In this section, we assess the feasibility of our modeling approach by calculating the percentage of misclassification for all $N=69$ subjects (including normal and AD). For each subject $k$, we denote $n_{\text{test}}^{(k)}$ represent the number of times they appeared in the test set across 10 resamples, and $n_{\text{error}}^{(k)}$ denote the number of times they were misclassified by the RF classifier. The subject-specific misclassification rate is then calculated as:
[…]
Figure 8 displays the relationship between the misclassification percentage and Cognitive scores (COGSTAT or MoCA). An overall trend observed in both plots is that subjects labeled as AD-aging (MCI and AD diagnoses) tend to exhibit higher misclassification percentages compared to the Normal-aging drivers. This is evident in the middle cognitive score ranges (COGSTAT: 580–650; MoCA: 12–20), where the misclassification is higher. It suggests that the middle zones correspond to a transition region where the classifier shows greater uncertainty, due to the overlapping feature space between early-stage impairment and normal aging. It underscores the anticipated challenge of reliably distinguishing subtle cognitive impairments/MCI from normal aging. On the other hand, subjects with low (COGSTAT $<550$ and MoCA $<8$) and high (COGSTAT $>700$ and MoCA $>20$) cognitive scores consistently yield lower misclassification rates. This behavior indicates that the classifier learned prominent feature representations for cognitively distinct individuals in these regions and it is able to classify the aging groups more confidently. Additionally, it reinforces the validity of using COGSTAT and MoCA as effective biomarkers in well-separated diagnostic states. Overall, these trends are consistent with the model’s performance as reported in Table 3. The agreement between global performance metrics and localized misclassification patterns supports the interpretation that the model’s predictive success is not random but rather driven by the prominent vision signals yielded from a certain cognitive state. Subjects with well-defined cognitive profiles— clear AD aging (Alzheimer/strong MCI) or normal aging are more reliably classified. However, uncertainty increases in transitional cognitive states, reflecting appropriate model sensitivity to the complexity of early-stage classification. This behavior supports the validity of our approach and suggests that further improvement may require incorporating auxiliary clinical signals to resolve the transition zone uncertainty. Overall, this subject-level metric captures individual classification stability and uncertainty at the subject level rather than aggregated classification accuracy alone.
@@ 4.3 Limitations and Future Works
The current setting evaluates only two types of roadways – freeway interchanges and interstates. Future work can include a broader set of driving scenarios (e.g., urban intersections, rural roads) to generalize findings across different traffic and environmental conditions. The driver-level separation was limited to 3-fold cross-validation with a fixed number of test subjects. Although this approach evaluates inter-subject generalization, the small fold count and limited test pool can lead to variance in performance estimation. Recruiting more diverse drivers and conducting longitudinal assessments would strengthen generalization. The use of in-vehicle video data raises potential concerns regarding privacy. Future implementations should address these with robust data management.
Another notable limitation is the relatively limited availability of high-quality in-vehicle video data per cognitive group. The model’s misclassification rates, particularly in the transitional region, are likely influenced by this data scarcity. We believe that acquiring a larger and higher-quality dataset with enriched video samples for both cognitively normal and AD groups will allow the model to learn more distinct and generalizable feature representations. This would be especially critical for improving classification accuracy in cases with confusing behavioral features.
5 Discussion
The results suggest that scenario-based analysis plays a significant role in video-based cognitive assessment. By segmenting driving video data into semantically distinct scenarios, our approach enables targeted analysis of driving behavior under specific cognitive demands. In the scope of this work, freeway interchange scenarios yield significantly higher classification performance compared to interstate scenarios under driver-level separation. Therefore, certain driving scenes inherently capture richer behavioral variance, which large vision models can leverage to distinguish between Normal vs. AD aging patterns. This approach offers a low-cost, scalable alternative to traditional cognitive screening methods, enabling passive and non-intrusive cognitive monitoring. It has direct implications for the healthcare industry, where early detection of cognitive decline remains a costly and time-consuming process. Moreover, the ability to extract digital biomarkers from routine driving behavior transforms personal vehicles into affordable diagnostic tools. This reduces patient burden while supporting continuous and longitudinal health assessment. The approach holds promise in providing an efficient digital health solution for aging populations in early detection of cognitive decline and screening in underserved or rural areas with limited access to healthcare facilities.
@@ 6 Acknowledgements
7 Appendix

### v2

@@ 3.2 Executing Video-based Model
We conducted the section 2.4 experiments across the two representative scenarios with both transitional and straight driving maneuvers. The aim was to investigate whether the large vision model could capture the subtle visual features to effectively distinguish driving behavior relevant for cognitive assessment. The experiments were conducted under the following sampling protocols (step 4 in section 2.4):
@@
    Random Sampling: Video segments are randomly split into train and test sets without enforcing driver-level separation.
@@
    Driver-level Separation (DLS): A leave-$k$-drivers-out cross-validation strategy is employed, where $k=5$ drivers were excluded during training and used exclusively for testing. This protocol was repeated over $r=3$ independent runs with different left-out driver groups and the final metrics were determined by averaging those runs.
In our experiments, the parameters were set as: (1) Input image size (960x752x3), (2) Vision model backbone: ResNet-v2 [41] (3) Video-to-vector dimension (6144x1), (4) Frame rate $\{1,10\}$ Hz, (5) Low-dimension size, $n\in\{50,100,200\}$, and (6) Distance metric: Euclidean distance. This setting provided a comprehensive understanding of how video-derived feature spaces behave under varying sampling conditions and data splits and laid the foundation for downstream classification tasks.
@@ 4.1 Model Performance on Different Scenarios
As discussed in section 2.4, we trained two RF models to identify the aging-groups (Normal-aging vs. MD-aging) across the Freeway interchange and Interstate scenarios. Table 3 reveals the performance of the trained RF models. The $\Delta$ row (without driver-level separation) indicates significant separation on all metrics $\mathbf{(a,P,R,F1)}$ and validates the selection of these contrastive scenarios to classify aging groups. The performance difference (grey rows) between the two scenarios indicates that there are prominent vision features in these specific scenarios that can successfully distinguish the aging groups solely by looking at the driving behavior features.
@@
Furthermore, the driver-level separation yields slightly widened separation, which can be observed on all metrics $\mathbf{(a,P,R,F1)}$ in the $\Delta_{\text{ds}}$ row. It surpassed the performance difference achieved without driver-level separation; for example, the accuracy difference went up from $\Delta$ = 15.91% to $\Delta_{\text{ds}}$ = 17.64%, indicating a 1.73% increment in performance difference. The $\Delta_{\text{ds}}$ aligns with achieving a significant performance difference while there is no “subject-level overlapping”. It also validates that the Fwy-int scenario includes more prominent driving movement features compared to its counterpart. On the other hand, the Interstate scenario performs significantly worse, yielding lower accuracy and F1-score. Therefore, the Freeway interchange is a more behaviorally informative and cognitively sensitive driving scenario, where complex maneuvers, short distances and lower speed variability have more nuanced driver responses.
@@
We compared our results with a recent study [42] that used a multi-modal (demographics and sleep log data) framework to detect cognitive decline in older drivers. Our method outperformed both without Driver-level Separation ([42] accuracy 68.64% vs. our Fwy-int. accuracy 71.03%) and with Driver-level Separation ([42] accuracy 70.48% vs. our Fwy-int. max accuracy 72.32%) cases on the RWRAD dataset. This highlights the effectiveness of scenario-specific vision features as non-invasive digital biomarkers for cognitive aging classification. Moreover, another recent study by Danda et al. [43] found that amyloid-positive older adults exhibited higher average speeds, intermittent speed variability and increased acceleration on freeway on-ramp driving (similar to Fwy-int scenario), suggesting less adaptive driving. In clinical research, Amyloid is a protein that accumulates abnormally in the brain and forms plaques linked to the development of MD. Patients with cognitive impairment who test amyloid-positive are typically diagnosed as having MD, whereas amyloid-negative patients are generally excluded from an MD diagnosis [44, 45]. These findings confirm that Freeway ramp driving requires increased cognitive load involving behavioral reactions to rapidly changing traffic and abrupt speed shifts, combining perception with motor control [43]. Additionally, they underscore our results of higher classification accuracy ($\mathbf{a}$ = 71.03%) and F1-score ($\mathbf{F1}$ = 0.6712) in Freeway interchange scenarios compared to the Interstate settings ($\mathbf{a}$ = 55.12%, $\mathbf{F1}$ = 0.5368). Although their approach combined driving patterns with physiological signals, it validates the relevance of using contrastive and challenging roadway scenarios in cognitive assessment tasks. Therefore, this study supports our findings that the Freeway interchange scenario provides more informative and strong behavioral cues that the large vision model can leverage to better separate cognitive aging groups.
@@ 4.2 Percentage of Misclassification
In this section, we assessed the feasibility of our Vision model-based approach by calculating the percentage of misclassification for all $N=69$ subjects (including normal and MD). For each subject $k$, we denote $n_{\text{test}}^{(k)}$ represent the number of times they appeared in the test set across 10 resamples, and $n_{\text{error}}^{(k)}$ denote the number of times they were misclassified by the RF classifier. The subject-specific misclassification rate is then calculated as:
% $\text{\% of missclassification}^{(k)}=100\times\left(\frac{n_{\text{error}}^{(k)}}{n_{\text{test}}^{(k)}}\right)$
Figure 8 displays the relationship between misclassification percentage and cognitive scores. An overall trend observed was that subjects labeled as MD-aging tend to exhibit higher misclassification percentages compared to the Normal-aging drivers. This is evident in the middle cognitive score ranges (COGSTAT: 580–650; MoCA: 12–20), where the misclassification is higher. It suggests that the middle zones correspond to a transition region where the classifier shows more uncertainty, due to the overlapping feature space between early-stage impairment and Normal-aging. In addition, it underscores the anticipated challenge of reliably distinguishing subtle cognitive impairments from Normal-aging. On the other hand, subjects with low (COGSTAT <550<550 and MoCA <8<8) and high (COGSTAT >700>700 and MoCA >20>20) cognitive scores consistently yield lower misclassification rates. This behavior indicates that the classifier learned prominent feature representations in these regions and it is able to classify the aging groups more confidently. Additionally, it reinforces the validity of using COGSTAT and MoCA as effective biomarkers in well-separated diagnostic states. Overall, these trends are consistent with the model’s performance as reported in Table 3. The global performance and localized misclassification patterns support the interpretation that the model’s predictive success is not random but rather driven by the prominent vision signals yielded from a certain cognitive state. Subjects with well-defined cognitive profiles were more reliably classified. However, uncertainty increased in transitional cognitive states, reflecting appropriate model sensitivity to the early-stage classification. This behavior supports the validity of our approach and suggests that further improvement may require incorporating auxiliary clinical signals to resolve the transition zone uncertainty. Overall, this subject-level metric captures individual classification stability and uncertainty at the subject level.
@@ 4.3 Limitations and Future Works
The current dataset is constrained to a limited number of driving scenarios, which might not cover the full range of real-world driving complexity (e.g., night driving, inclement weather, heavy traffic). Future work can include a broader set of driving scenarios to generalize findings across real-world traffic and environmental conditions. The driver-level separation was limited to a 3-fold cross-validation with a small pool of test subjects. This can lead to variance in performance estimation and may not generalize well to other populations and road infrastructures. Recruiting more diverse drivers and conducting longitudinal assessments would strengthen generalization. Additionally, the number of dementia subjects in the study was relatively small ($N=9$). Including more dementia participants would guarantee the statistical validity of the findings and provide deeper insights into the driving cues associated with dementia. Moreover, the use of in-vehicle video recording raises potential concerns regarding privacy. Future implementations should address these with robust data management. Another notable limitation is the relatively limited availability of high-quality in-vehicle video data per cognitive group. The model’s misclassification rates, particularly in the transitional region, were influenced by this data scarcity. We believe that scaling up the dataset with higher-quality video samples for both cognitively normal and dementia groups will allow the method to extract more meaningful and generalizable features. This would be especially critical for improving video analysis accuracy in cases with confusing behavioral features.
5 Conclusion
This study highlights the potential of monitoring real-world driving behavior as a valuable tool for early cognitive health assessment in older adults. The results suggest that scenario-based analysis plays a significant role in video-based cognitive assessment. By segmenting driving video data into semantically distinct scenarios, our approach enables targeted driving behavior analysis under specific cognitive demands. In the scope of this work, freeway interchange scenarios yield significantly higher classification performance compared to interstate scenarios under driver-level separation. Therefore, certain driving scenes inherently capture richer behavioral variance, which large vision models can leverage to distinguish between cognitive aging driving patterns. This approach offers a low-cost, scalable alternative to traditional cognitive screening methods, enabling passive and non-intrusive cognitive monitoring. It has direct implications for the healthcare industry, where early detection of cognitive decline remains a costly and time-consuming process. Additionally, the ability to extract digital biomarkers from routine driving behavior transforms personal vehicles into affordable diagnostic tools. This reduces patient burden while supporting continuous and longitudinal health assessment. Moreover, it holds promise in providing an efficient screening tool for aging populations in early detection of cognitive decline, especially in areas with limited access to healthcare facilities.
@@ 6 Acknowledgements
Appendix A Study Recruitment Criteria
The inclusion and exclusion criteria used for participant recruitment are summarized in Table 4.
Table 4: Participant inclusion and exclusion criteria [32]
Inclusion Criteria Exclusion Criteria
(a) age between 65 and 90 years (a) pulmonary disease requiring chronic medication
(b) residing independently in Nebraska (b) Congestive heart failure
or nearby states (e.g., Iowa) (c) major psychiatric illness
(c) possession of a valid driver’s license (d) active vestibular disease
(d) met Nebraska visual acuity standards (>20/50>20/50 OU, corrected or uncorrected) (e) substance abuse within one year prior to enrollment
(e) proficiency in English (f) use of medications that could confound study outcome
(f) capability to provide informed consent (g) clinical or neurological conditions (e.g., stroke) that could interfere with the study outcome
Appendix B Cognitive Tests used for COGSTAT score
Cognitive Tests used for calculating COGSTAT score in the RWRAD dataset across five cognitive domains [9] are summarized in Table 5.
Table 5: Cognitive Tests conducted for the COGSTAT composite score.
Cognitive Domain Cognitive Test
(1) Memory Benson, Recall
Craft Story Delay, Verbatim
HVLT, Delay
(2) Language Category Fluency, Animals
Category Fluency, Vegetables
MINT
(3) Visuospatial Benson, Copy
WAIS Blocks
(4) Executive Function Trails B
Verbal Fluency (F)
Verbal Fluency (L)
(5) Attention Trails A
Number Span Forward (Total correct)
Number Span Backward (Total correct)
Note: If all cognitive tests are the same as the mean, final COGSTAT will be $14\times 50=700$